OpenStudy (anonymous):

which quadratic function, when graphed, has x-intercepts of 4 and -3?

OpenStudy (anonymous):

An equation with zeros at 4 and -3. ie. \[(x-4)(x+3)\] If you expand that; \[x^2-4x+3x-12=x^2-x-12\]

OpenStudy (anonymous):

a. y= (x=3)(x+4) b. y=(x+3)(2x-8) c. y= (3x-1)(4x+1) d. y= (3x+1)(8x-2)

OpenStudy (anonymous):

Find the zeros of all those. The equation I gave above was one of the infinite possible forms. Any equation with zeros of 4 and -3 would be valid. I would take a look at b.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem!